OpenStudy (lena772):

Part A: Maria rented a coat at $285 for 3 days. If she rents the same coat for 6 days, she has to pay a total rent of $510. Write an equation in the standard form to represent the total rent (y) that Maria has to pay for renting the coat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (lena772):

If she rents for 3 days I find that she pays $95 per day, however if she rents for 6 days I find that she pays $85 per day

OpenStudy (lena772):

so i'm stuck now

OpenStudy (lena772):

@hartnn

OpenStudy (the_fizicx99):

Your x,y values for part A is?

OpenStudy (the_fizicx99):

3,285 6,510 So f(b) - f(a) / b - a

OpenStudy (the_fizicx99):

Plug in, i gotta get off :p

OpenStudy (lena772):

f(6)-f(3)/6-3

OpenStudy (lena772):

510-285/6-3 225/3 75

OpenStudy (the_fizicx99):

Yes. lol I might make typo's.. On my phone

OpenStudy (lena772):

so that's my equation for A?

OpenStudy (the_fizicx99):

Y= 75x + ? What do you think b is?

OpenStudy (lena772):

y=75x+b 510=75*6+b 510=450+b 60=b

OpenStudy (lena772):

y=75x+60

OpenStudy (lena772):

B is f(x)=75x+60

OpenStudy (lena772):

i have no idea how to do C

OpenStudy (the_fizicx99):

K so why'd you put 60 as b?

OpenStudy (the_fizicx99):

Ohh you put 6 as # of days she rented, right?

OpenStudy (lena772):

:/ What?

OpenStudy (the_fizicx99):

I don't get how you got 60 as b, could you explain

OpenStudy (lena772):

I calculated the y intercept

OpenStudy (lena772):

y=75x+b replace y with a y value from the chart=75*(replace x with the corresponding x value)+b 510=75*6+b 510=450+b 60=b

OpenStudy (the_fizicx99):

Okay, for C. Part C: I would graph the days on the x axis starting at 0 through 10 with intervals of 1. Then graph the cost on the y axis starting at 0 through 510 with intervals of ? What do you think "?" is

OpenStudy (lena772):

what "?"

OpenStudy (the_fizicx99):

What intervals would you use? "Purpose is to guide learners " -OS .. Sorry if this seems annoying :/

OpenStudy (lena772):

Um I'd put days on the x axis because time is an independent variable and i'd put the integers from 0-10 because it's only logical that you have integers when counting days

OpenStudy (lena772):

for the y axis I'd go up in integers of 25 or 50

OpenStudy (lena772):

intervals*

OpenStudy (the_fizicx99):

Yeah I explained the x-axis, I mean what intervals would u use for the y... Would you agree with 75? xD

OpenStudy (the_fizicx99):

@Lena772 get it?

OpenStudy (lena772):

yes i'd agree with 75

OpenStudy (the_fizicx99):

So does that answer your question?

OpenStudy (lena772):

yes thanks

OpenStudy (lena772):

Check your messages >.>

OpenStudy (the_fizicx99):

I did <.<